

********************************************************************
UARS WEEKLY STATUS REPORT - 2 July 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring scans
for periods of UARS real-time telemetry or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2002079. MEPS data files between
2002080 and 2002119 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of April 6-7 of 2000, March 31 - April 1 of 2001, April 11-12
of 2001, and November 6-7 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 19 June 2004 (Orbit 69886, GMT Day 171) through
25 June 2004 (Orbit 69991, GMT Day 177).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 171/2259 69901 POWER OFF
PEM/AXIS 174/2250 69946 POWER OFF
PEM 177/1221 69984 POWER OFF (S/C LOW PWR)
The beta angle ranged between the angles of 52.0 degrees to 76.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 27 June 2004 (DOY 180).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.26 to 1.27 degrees C. Additional
battery performance data for each battery on 18 June and 25 June 2004
are provided in Appendix D.
The clock error ranged from 11.5 to -12.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
HALOE = 01
SUSIM = 02
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week UARS is flying through its high beta period. The power
margin during this period is always very difficult to estimate. The
FOT will be closely monitoring power and adjusting instrument status
throughout this period if needed. Instrument teams will be informed
if the status of their instrument is changed.
This week the FOT received the new Eclipse software build. It has
been installed on the prototype machine and testing will begin next
week. There were several DR fixes incorporated in this build, so
testing will take a few weeks before this build will be ready for
operations.
MEETINGS
The FOT attended the biweekly meeting with the HRDI team to discuss
the attitude error seen in their data. FDF and Joe Hashmall continue
to perform independent analysis to try to isolate the source of this
problem. Their focus for the next few weeks will be the TAM residuals
and digitization affects and the UFLTR tuning.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 25 June 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69898 171/1857 04-172 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
69901 171/2259 04-173 HALOE POWER OFF
69911 172/1459 04-174 SUSIM EXCEPTION CLEAR
69925 173/1413 04-175 SUSIM AZIMUTH COMMANDS
69931 173/2254 04-176 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69946 174/2250 04-177 PEM/AXIS POWER OFF
69950 175/0503 04-178 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
69951 175/0744 04-179 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
69980 177/0459 04-180 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
69984 177/1221 04-181 PEM POWER OFF (S/C LOW PWR)
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 25 June 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 June 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
69898 171/1857 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
69931 173/2305 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69950 175/0503 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
69951 175/0748 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
69980 177/0459 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 June 2004
18 June 2004 - GMT Day 170 Beta = 46.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-179.2 +4.83/+3.87 +25.2/-14.4 25.4 1.03 22.0
3 +16.8/-112.0 +3.87/+2.60 +24.8/-14.4 25.4 1.02 21.0
25 June 2004 - GMT Day 177 Beta = 76.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 +05.6/-156.8 +2.60/+1.34 +14.4/-14.4 24.3 1.01 12.6
3 +11.2/-089.6 +1.97/+0.71 +14.4/-14.4 24.3 1.00 12.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 25 June 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT 69952 175/0927 30832 Generic Late Acquisition
04-032 FOT 69981 177/0747 30832 Generic Late Acquisition
04-032 FOT 69988 177/1759 30832 Generic Late Acquisition
04-032 FOT 69989 177/1949 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-033 FOT 69861 169/0702 36148 43 Second Telemetry
Dropout
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT See List June 04 30832 Generic Late Acquisition